[Brahms-l] Shift report 20040330 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Tue Mar 30 2004 - 00:03:08 EST
------------------------------------------------------------
Date: Tue Mar 30 00:03:07 2004, Shift: 16:00-24:00
Supervisor: JHL, with: 
Log book#15 pages 102 - 102

Shift Summary:
------------------------------------------------------------
1. Beam for 5 hrs and 20 min
   -1 full store staring at ZDC~650/sec
   -45 bunches after PHOBOS field flip
   -background was at high-side and varing 
    tofw/zdc ~ 1.5 -> 4 -> 2 -> 3 -> 2 over the store
   - currently no beam: "Controls failure"   

2. Data
   -MRS 45 deg 6B 1M program done
    Another 1M trigger3 to go at 6B
   -FS 6deg 1/4A 1M program done
    Now 6deg 1/4B 1M program (350K trigger6 so far)

3. No detector problems 
 
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Mar 30 00:03:40 2004

This archive was generated by hypermail 2.1.8 : Tue Mar 30 2004 - 00:04:11 EST